OpenStudy (anonymous):

Calculate the slope of the tangent to each curve at the given value of x using first principles. f(x)= (x+4)^1/2 , x=5

OpenStudy (anonymous):

first principles? The easiest way to do this is with calculus, by taking the derivative of f(x). \[f'(x)=1/2(x+4){^{-1/2}}\] and then plugging in the x=5, so f'(x)=-1/6 and 1/6

OpenStudy (anonymous):

Thank you for responding to my question. I know how to do the question using derivates but I can't quite remember how to do it with first principles because we'll be asked to do so on the exam.

OpenStudy (anonymous):

Hey, so I'm not sure what first principles are. Could you explain those?

OpenStudy (anonymous):

m= f(x+h) - f(x) / h basically the harder way to find the slope than using derivatives. lol

OpenStudy (anonymous):

Ok I think I got them So m= f(x+h) - f(x) / h So then the slope is \[m=Deltay/Deltax=dy/dx\] So \[dy/dx=f(x+h)-f(x)/h\] Thus, \[dy/dx=\lim_{h \rightarrow 0} f(x+h)-f(x)/h\]\] So since f(x)=(x+4)^1/2, f(x+h)=(x+4+h)^1/2 So the equation is \[m=(x+4+h) ^{1/2}-(x+4)^{1/2}/h\] \[dy/dx=\lim_{h \rightarrow 0}(x+4+h) ^{1/2}-(x+4)^{1/2}/h\]

OpenStudy (anonymous):

oh right okay, thank you!

OpenStudy (anonymous):

I don't know how to take the limit of that, but at least it's a start :)

OpenStudy (anonymous):

yep I know where to go from there so thanks :)